Prev Next

SW&IT Meeting Minutes 2018-08-27

Attendees: Andrew, David, Igor, Kellee, Matthieu, Petr, Yang

Outlook

News from management

  • staff announcement
  • SSD lessons learnt

Previous Action Items

  • Vsphere (DE)
  • power plan update draft (DE)
  • IT email format (DE) DONE
  • IT email remaining dup (DE)
  • LBTI/jumbo switch (DE) DONE on-order

Open topics

Bring any topics you want for discussion at the meeting here. Use <topic>[<name>] formatting.

  • trivial passwords (ovms-uei, others?) (SH when he comes back)
  • OVMS racktangle upgrade checkout + announcement
  • AGW2 umac comm, issues
  • KS tech-talks
  • TCS release planning (PK)

This week in github

Team reports

YZ

  • Working on pointing and collimation documentation.
  • Working on experimenting an LBTO Python 3 module and collections, as well as its documentation.
  • Will work with Petr on evaluating new collimation tools, and then with John Hill perhaps?
  • Met with Al Conrad on new AO project. Just getting started and will focus on overall design and some background research at the moment.

MB

  • management OC, (no WPC)
  • pp2018 completion, pp2019 planning
  • AO tools meeting and moving forward
  • PP2019 meeting
  • OVMS+TT meeting
  • PMC support
  • LBC skytech meeting w. Fernando
  • SHARK-VIS meeting w. Fernando
  • Labor reporting
  • AGW2 UMAC comm. issues
  • wiki software cleanup

next:
  • IT to github issues
  • AC/KS transition planning
  • science archive follow up

KS

Last week
  • Recovery that was overlooked during the power down/up:
    • weatherstation software wasn't running, no AllSky
    • EPICS IOCs were down, so Status Server and Alarm Handlers were incorrect
    • ovms and dimm computers needed reboot for mounts
    • tried to clear up some confusion about how the startup/shutdown script works in regard to manual vs automatic (it cannot do anything automatically if it's commented out in the cron table)
  • OVMS is working with 12 boards now - John reseated board 8 that was showing up as "empty" in the ping test program. However, we lost the application when power went down or it was rebooted. Petr and LeRoy fixed this write problem with the SD card.
  • Committed and tagged OVMS updates for new kernel. Had Elwood check out the LBTI software and he verified he could still connect and get reasonable data.
  • Went through my notes for the LBC firmware upgrade to dataprobes. Turns out I cannot do it remotely.
  • Re-acqainted myself with the LBC controller upgrade and where we left things. Attended a meeting to get status from Fernando. Sent Fernando the last list of issues I had and the wiki pages I kept for testing.
  • Updated the lbc.zip release with new config files from Fernando's detector tuning on the mountain.

This week
  • Release 2018A patch for GCS for OAC upgrade. Test on the mountain with AGW4.
  • Update TCS master branch for 2018B with GCS updates for OAC HK.
  • Fix build of OAC clients to not include telemetry library.
  • Document the OVMS kernel upgrade.

AC

Last week
  • IT #7368 - required running processes/dependencies for ARGOS telemetry data
  • Acetri and ARGOS Wiki conversion to Foswiki
  • Migrate TBAD and LCH into Git
  • Build/install ARGOS and TAN with tests in Jenkins

This week
  • Continuation of the above
  • Provide ARGOS docker image to MK (MPIA)
  • Responsibility transition with KS

PK

Last week
  • 2018B release (ready on tcs-test)
  • C++11 lib-telemetry, running tests
  • OSS patch for 2018A (DX M3 home offset change)

This week
  • SX M1 crate help (M1 not running yet, both crates are downtown)
  • 2018B (expecting to switch 4th September)

IS

This Week
  • LUCI
    • LUCI RTD
    • Possibly other LUCI AFC testing if Dave gets to test AFC
    • HandOver document, (virtually done, need to find three Doug Summers documents and load them to the VCAN)
  • PIT
    • Michelle's team will test the PIT and give me and let me know of any issues.
Next Week
  • TBD but most probably will be LUCI centric
Topic revision: r10 - 10 Dec 2018, UnknownUser
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback